A bottle of a commercial reagent-grade hydrochloric acid, HCl has the following information on its label: Assay 37% HCl by mass; density 1.19 g/mL. Calculate the volume (mL) of concentrated HCl needed to prepare 250 mL of 0.1M HCl.


1
Expert's answer
2021-01-22T06:22:56-0500

To find the molar concentration of concentrated acid, it is reasonable to assume the volume of 1000 mL (1 L) of it, and to calculate the number of moles of HCl in this solution.

The mass of concentrated HCl:

"m=D\times{V}=1.19g/mL\times1000mL=1190g"

The mass of pure HCl in the concentrated acid:

"m=1190g\times0.37=440.3g"

Amount of HCl in moles:

"n=\frac{440.3g}{36.46g/mol}=12.1mol"

Therefore, the concentration is 12.1M


Now, knowing the concentration of both solutions and the volume of the diluted one, the volume of the concentrated HCl can be calculated using the formula "c_1V_1=c_2V_2". Therefore,


"V_1=\frac{c_2V_2}{c_1}=\frac{250mL\times0.1M}{12.1M}=2.07mL"


Answer: 2.07 mL